# Consistent energy-momentum trace couplings of fluids

### v1

@@ 2.4 Thermodynamic quantities
We now revisit the equations of motion Eqs. (2.29) and (2.30) using a thermodynamic approach. The first law of thermodynamics can be expressed locally by
@@
We see that the effect of the non-minimal trace coupling is a temperature contribution given by
@@
$(\upmu-\upmu_{\rm tr})U_{\mu}+\varphi_{,\mu}+s\theta_{,\mu}+\beta_{A}\alpha^{A}_{,\mu}=0\,.$ (2.40)
where $\upmu_{\rm tr}$ is the contribution of the non-minimal trace coupling to the chemical potential given by
@@
The modifications of Eq. (2.29) and Eq. (2.30) are due to matter tensor couplings which contribute to the temperature and chemical potential. Since the matter tensor of a fluid will always depend on the particle number density, the term $\upmu_{\rm tr}$ will in general contribute to the fluid’s equations of motion.
@@ 2.5 Conservation equations via field equations
We now check the conservation equations for the effective energy-momentum tensor. Typically, the continuity equation gives $\nabla^{\mu}T^{\rm eff}_{\mu\nu}=0$. This is not true in general in the context of the coupling with between geometry and matter trace. We can easily see this by taking the divergence of the field equations, Eq. (2.20)
@@
While this conservation equation is valid in general, it will be convenient for us to consider its projection along the fluid’s 4-velocity and along the perpendicular directions. To do this, we define the usual projection $h^{\mu\nu}=U^{\mu}U^{\nu}+g^{\mu\nu}$.
First, contracting with $U^{\nu}$
@@
where we have used the relations $U^{\mu}h_{\mu\nu}=0$, and $U^{\nu}\nabla^{\mu}U_{\nu}=0$. Noticing also that $U_{\mu}\nabla^{\mu}s=0$ and $U_{\mu}\nabla^{\mu}n=-n\nabla^{\mu}U_{\mu}$, we have
$2\kappa U^{\nu}\nabla^{\mu}T_{\mu\nu}$ $=-\Bigl(f_{T}n\frac{\partial T}{\partial n}\Bigr)\nabla^{\mu}U_{\mu}+f_{T}\Bigl(\frac{\partial T}{\partial n}n\nabla^{\mu}U^{\mu}\Bigr)=0\,.$ (2.48)
@@ 2.6 Conservation equations via fluid equation
We will now re-derive the the conservation equation from the fluid equation using only the equations of motion, without referring to the field equations. This will further verify the internal consistency of our equations and can be seen as the final step in demonstrating the correctness of our approach.
By applying the covariant derivative to the energy-momentum tensor one simply gets $\nabla_{\mu}T^{\mu\nu}=\nabla_{\mu}(\rho U^{\mu}U\nu)+\nabla_{\mu}(ph_{\mu\nu})$. We will evaluate this along the flow and in the perpendicular direction. If we contract this equation with $U_{\nu}$, it is straightforward to verify that $U_{\nu}\nabla_{\mu}T^{\mu\nu}=0$ which matches the result of Eq. (2.48).
To evaluate $h_{\nu\sigma}\nabla_{\mu}T^{\mu\nu}$, we begin with [7, Eq. (2.13)]
@@
Using the fluid’s equations of motion removes two out of the three final terms
@@
which can also be found in [7] and is unaffected by the trace coupling. Finally, one uses Eq. (2.29)
@@
Next, we can expand the skew-symmetrisation and perform the covariant differentiation using the product rule, which gives three terms. These can be slightly rewritten and combined in such a way that one arrives at
@@ 3 Schutz’s velocity potentials approach
In this section, we work with the relativistic perfect fluid Lagrangian given by Schutz [8]. We follow the structure of the previous section to compare between the two pictures and verify our approach of using the off-shell Lagrangian. Here, the Lagrangian is given simply by $\mathcal{L}_{\rm fluid}=p\,\sqrt{-g}$. The full description of the fluid is contained in the velocity potentials. The 4-velocity unit vector is defined as a function of the independent variables with $U_{\nu}=\upmu^{-1}(\phi_{,\nu}+\alpha\beta_{,\nu}+\theta s_{,\nu})=\upmu^{-1}V_{\nu}$. The fluid action for the relativistic perfect fluid in this approach is
@@
with $\upmu=\sqrt{-g^{\mu\nu}V_{\mu}V_{\nu}}$. Thus, while $\upmu$ is a potential in this approach, it is not an independent variable. The independent variables which we vary the action with respect to are:
@@
We already notice here the analogue between $\rho$ and $p$ in Brown’s and Schutz’s formalisms, a feature that will continue to emerge throughout the section. In a similar fashion, we start by deriving the field equations and equations of motion, show the on-shell reduction, thermodynamic quantities, and conservation equations.
@@ 3.1 Gravitational field equations
We start again by varying the action for $f(R,T)$ model, this time with a matter Lagrangian density given by Eq. (3.1). We use our previous results in Section 2.1 since the difference in this picture is only in the trace of the fluid matter and the Lagrangian. We have
@@
The variation of the energy-momentum trace $\delta T$ can be derived from Eq. (3.3)
@@
We substitute the variation $\delta\upmu$ back in $\delta T$ to get the final expression in terms of independent variables
@@
Now we collect all the terms and write the action variation in terms of the independent variables
@@
Notice that here the extra term couples to 4-velocity unit vectors $U_{\mu}U_{\nu}$ but not the metric, which, by using the definition of $T_{\mu\nu}$, suggests an effective energy density rather than pressure, given by
@@
Not surprisingly, we get analogous results to that found in the Brown picture. Here, we see the exchange of the derived quantity between $p$ in Brown’s picture and $\rho$ in Schutz’s picture led to the exchange of their effective role in the field equations. The expression for $\rho_{\rm eff}$ agrees with the on-shell expression in the argument presented in Section 2.3. Of course, one has to be careful when changing the variables $n$ and $\upmu$, however, it should be clear at this point that the trace coupling can only yield a contribution that affects the energy density or the pressure.
@@ 3.2 Equations of motion
$n\equiv\frac{\partial p}{\partial\upmu}+\frac{f_{T}}{2\kappa}\frac{\partial T}{\partial\upmu}\,$, (3.19)
then we find from Eq. (3.17) that $\partial_{\nu}\big(nU^{\nu}\big)=0$. Using this relation and the product rule in Eqs. (3.15) and (3.18) allows us to write the equations of motion as:
@@
The equations of motion (3.20)–(3.23) now match the expected result found by Schutz [8] with the new definition of $n$ in Eq. (3.19). We see, however, that the entropy variation $\delta s$ leads to an additional term $(2\kappa)^{-1}f_{T}\partial T/\partial s$ which is a result of the non-minimal coupling between curvature and the matter trace. In fact, even the factor $n$ in Eq. (3.23), which ensures the conservation of the particle number now depends on the non-minimal coupling between $R$ and $T$. Those effects will be discussed in more detail in Section 3.4 and Section 4.
@@ 3.3 On-shell fluid Lagrangian
$S_{\textrm{fluid}}=\int p(\mu,s)\,\sqrt{-g}\,d^{4}x\qquad\textrm{off-shell}\,$, (3.25)
@@
$S_{\textrm{fluid}}$ $=-\int\rho_{\textrm{eff}}\,\sqrt{-g}\,d^{4}x\qquad\textrm{on-shell}\,.$ (3.31)
Therefore on-shell, we do get back the expected expression after utilizing the equations of motion. This not only ensures the consistency of our results so far, but also strengthens the argument that one indeed has to be careful when working with the on-shell expression of the Lagrangian as it only holds up when the equations of motion are considered. However, when applying a variation principle, one must resort to the off-shell Lagrangian to get the full expression Eq. (3.10).
@@ 3.5 Conservation equations via field equations
We again check the consistency of our approach by first deriving the conservation equations from the field equations in this section and then derive them once more from the fluid equation in the following section.
@@
where the contracted derivative of the entropy term is eliminated by the equation of motion Eq. (3.22). By substituting the expression for $T_{\mu\nu}^{\rm eff}$ from Eq. (3.12) we get:
@@
Using the identities $h^{\nu}_{\sigma}h_{\mu\nu}=h^{\nu}_{\sigma}g_{\mu\nu}$ and $h^{\nu}_{\sigma}\nabla^{\mu}h_{\mu\nu}=h^{\nu}_{\sigma}U_{\mu}\nabla^{\mu}U_{\nu}$ allows us to combine the first two terms. After multiplying by $2\kappa$, the expression can be written in a form similar to that in Eq. (2.51)
@@
We have obtained the two expressions for the conservation equation parallel to the flow, Eq. (3.39) and perpendicular to the flow, Eq. (3.42). Now, we attempt to re-derive the same expressions using the fluid equation without resorting to the field equations.
@@ 3.6 Conservation equations via fluid equations
$\nabla^{\mu}T_{\mu\nu}$ $=\nabla^{\mu}\Big(\upmu\frac{\partial p}{\partial\upmu}U_{\mu}U_{\nu}\Big)+g_{\mu\nu}\nabla^{\mu}p\,.$ (3.43)
@@
By expanding the derivative of the pressure and using Eq. (3.22)
@@
which matches the result from the field equations in Eq. (3.39), verifying our previous result that in this picture, the parallel contraction of the energy-momentum divergence does not vanish!
Now, we contract with $h^{\nu}_{\sigma}$ to check the conservation equations in the perpendicular direction
@@
The term in round brackets vanishes (check Eq. (A.44) in Appendix A.5), leaving us with
@@
With that we have strong confidence that the correct approach to vary the fluid action is to use the full off-shell fluid Lagrangian. We showed that this can be done in both Brown’s approach with Lagrange multipliers or Schut’z approach based on velocity potentials. The results in both cases are consistent and lead to the same conclusions.

### v2

@@ 2.4 Thermodynamic quantities
We now revisit the equations of motion (2.29) and (2.30) using a thermodynamic approach. The first law of thermodynamics can be expressed locally by
@@
We see that the effect of the nonminimal trace coupling is a temperature contribution given by
@@
$(\upmu-\upmu_{\rm tr})U_{\mu}+\varphi_{,\mu}+s\theta_{,\mu}+\beta_{A}\alpha^{A}_{,\mu}=0\,$, (2.40)
where $\upmu_{\rm tr}$ is the contribution of the nonminimal trace coupling to the chemical potential given by
@@
The modifications of Eqs. (2.29) and (2.30) are due to matter tensor couplings which contribute to the temperature and chemical potential. Since the matter tensor of a fluid will always depend on the particle-number density, the term $\upmu_{\rm tr}$ will in general contribute to the fluid’s equations of motion.
@@ 2.5 Conservation equations via field equations
We now check the conservation equations for the effective energy-momentum tensor. Typically, the continuity equation gives $\nabla^{\mu}T^{\rm eff}_{\mu\nu}=0$. This is not true in general in the context of the coupling with between geometry and matter trace. We can easily see this by taking the divergence of the field equations (2.20):
@@
While this conservation equation is valid in general, it will be convenient for us to consider its projection along the fluid’s four-velocity and along the perpendicular directions. To do this, we define the usual projection $h^{\mu\nu}=U^{\mu}U^{\nu}+g^{\mu\nu}$.
First, contracting with $U^{\nu}$,
@@
where we have used the relations $U^{\mu}h_{\mu\nu}=0$ and $U^{\nu}\nabla^{\mu}U_{\nu}=0$. Noticing also that $U_{\mu}\nabla^{\mu}s=0$ and $U_{\mu}\nabla^{\mu}n=-n\nabla^{\mu}U_{\mu}$, we have
$2\kappa U^{\nu}\nabla^{\mu}T_{\mu\nu}$ $=-\Bigl(f_{T}n\frac{\partial T}{\partial n}\Bigr)\nabla^{\mu}U_{\mu}+f_{T}\Bigl(\frac{\partial T}{\partial n}n\nabla^{\mu}U_{\mu}\Bigr)=0\,.$ (2.48)
@@ 2.6 Conservation equations via fluid equation
We will now rederive the the conservation equation from the fluid equation using only the equations of motion, without referring to the field equations. This will further verify the internal consistency of our equations and can be seen as the final step in demonstrating the correctness of our approach.
By applying the covariant derivative to the energy-momentum tensor one simply gets $\nabla_{\mu}T^{\mu\nu}=\nabla_{\mu}(\rho U^{\mu}U^{\nu})+\nabla_{\mu}(ph^{\mu\nu})$. We will evaluate this along the flow and in the perpendicular direction. If we contract this equation with $U_{\nu}$, it is straightforward to verify that $U_{\nu}\nabla_{\mu}T^{\mu\nu}=0$ which matches the result of Eq. (2.48).
To evaluate $h_{\nu\sigma}\nabla_{\mu}T^{\mu\nu}$, we begin with [see Eq. (2.13) in [11]]
@@
Using the fluid’s equations of motion removes two out of the three final terms:
@@
which can also be found in [11] and is unaffected by the trace coupling. Finally, one uses Eq. (2.29):
@@
Next, we can expand the skew-symmetrization and perform the covariant differentiation using the product rule, which gives three terms. These can be slightly rewritten and combined in such a way that one arrives at
@@
2.7 Effective equations and dark sector interactions
The nonvanishing right-hand sides of the conservation equations (2.51) and (2.6) demonstrate the nonconservation of the energy-momentum tensor of the fluid, a common feature of theories with matter-curvature couplings. One can study the effects of such additional terms in the context of dark sector interactions by identifying suitable energy-momentum tensor contributions. To arrive at effective field equations, we begin with
$f(R,T)$ $=R+\bar{f}(R,T)\,$, (2.60)
with $f_{T}=\bar{f}_{T}$ and $f_{R}=1+\bar{f}_{R}$. This choice of function will isolate the Einstein tensor. The field equations (2.20) can now be written in a form similar to the Einstein field equations:
$G_{\mu\nu}=\kappa T_{\mu\nu}-\frac{1}{2}\bar{f}_{T}n\frac{\partial T}{\partial n}(g_{\mu\nu}+U_{\mu}U_{\nu})+\frac{1}{2}\bar{f}g_{\mu\nu}-\bar{f}_{R}R_{\mu\nu}-\big(g_{\mu\nu}\Box-\nabla_{\mu}\nabla_{\nu}\big)\bar{f}_{R}\,$, (2.61)
with all terms due to the nonminimal coupling moved onto the matter side of the equation. The first term on the right-hand side is the fluid’s matter energy-momentum tensor. The remaining terms are matter and geometry dependent since $\bar{f}$ depends on the Ricci scalar and the matter tensor’s trace. These remaining terms can be associated with dark energy and dark matter. We note, however, that this split is not unique. We set the dark energy part to be
$T_{\mu\nu}^{\rm DE}:=\frac{1}{2\kappa}\Bigl(\bar{f}-\bar{f}_{T}n\frac{\partial T}{\partial n}\Bigr)g_{\mu\nu}\,.$ (2.62)
This is motivated by the form of the cosmological constant term $\Lambda g_{\mu\nu}$. Assuming that dark matter is pressureless, it is reasonable to define the following dark matter part:
$T_{\mu\nu}^{\rm DM}:=-\frac{1}{2\kappa}\bar{f}_{T}n\frac{\partial T}{\partial n}U_{\mu}U_{\nu}\,.$ (2.63)
The remainder of Eq. (2.61) is an additional matter component:
$T_{\mu\nu}^{\bar{f}}:=-\frac{1}{\kappa}\bigl(R_{\mu\nu}+g_{\mu\nu}\Box-\nabla_{\mu}\nabla_{\nu}\bigr)\bar{f}_{R}\,.$ (2.64)
Using this notation, we can rewrite the field equations in a convenient form:
$G_{\mu\nu}=\kappa\Bigl(T_{\mu\nu}+T_{\mu\nu}^{\rm DM}+T_{\mu\nu}^{\rm DE}+T_{\mu\nu}^{\bar{f}}\Bigr)\,.$ (2.65)
One can immediately show that
$\nabla^{\mu}T_{\mu\nu}^{\bar{f}}=-\frac{1}{2\kappa}\bar{f}_{R}\nabla_{\nu}R=:-Q_{\nu}\,$, (2.66)
which can be interpreted as an interaction vector $Q_{\nu}$. For the other components, one cannot derive an independent nonconservation equation that only depends on that same interaction. Therefore, using the above notation, we can state that
$\nabla^{\mu}\Bigl(T_{\mu\nu}+T_{\mu\nu}^{\rm DM}+T_{\mu\nu}^{\rm DE}\Bigr)=-\nabla^{\mu}T_{\mu\nu}^{\bar{f}}=Q_{\nu}\,.$ (2.67)
Consequently, we have an interaction between the fluid, the identified dark energy and dark matter components. However, one cannot easily separate this interaction into the more standard form where $\nabla^{\mu}T_{\mu\nu}^{\rm DM}=-\nabla^{\mu}T_{\mu\nu}^{\rm DE}$, as discussed, for example in [13, 14]. This follows from our fluid equations (2.25)-(2.28) that contain the geometrical couplings, which imply that $\nabla^{\mu}T_{\mu\nu}\neq 0$. Even absorbing $T_{\mu\nu}^{\bar{f}}$ into either dark energy or dark matter will not change this outcome.
@@ 3 SCHUTZ’S VELOCITY POTENTIALS APPROACH
In this section, we work with the relativistic perfect fluid Lagrangian given by Schutz [12]. We follow the structure of the previous section to compare between the two pictures and verify our approach of using the off-shell Lagrangian. Here, the Lagrangian is given simply by $\mathcal{L}_{\rm fluid}=p\sqrt{-g}$. The full description of the fluid is contained in the velocity potentials. The four-velocity unit vector is defined as a function of the independent variables with $U_{\nu}=\upmu^{-1}(\phi_{,\nu}+\alpha\beta_{,\nu}+\theta s_{,\nu})=\upmu^{-1}V_{\nu}$. The fluid action for the relativistic perfect fluid in this approach is
@@
with $\upmu=\sqrt{-g^{\mu\nu}V_{\mu}V_{\nu}}$. Thus, while $\upmu$ is a potential in this approach, it is not an independent variable. The independent variables which we vary the action with respect to are
@@
We already notice here the analog between $\rho$ and $p$ in Brown’s and Schutz’s formalisms, a feature that will continue to emerge throughout the section. In a similar fashion, we start by deriving the field equations and equations of motion and show the on-shell reduction, thermodynamic quantities, and conservation equations.
@@ 3.1 Gravitational field equations
We start again by varying the action for $f(R,T)$ model, this time with a matter Lagrangian density given by Eq. (3.1). We use our previous results in Sec. 2.1 since the difference in this picture is only in the trace of the fluid matter and the Lagrangian. We have
@@
The variation of the energy-momentum trace $\delta T$ can be derived from Eq. (3.3):
@@
We substitute the variation $\delta\upmu$ back in $\delta T$ to get the final expression in terms of independent variables:
@@
Now we collect all the terms and write the action variation in terms of the independent variables:
@@
Notice that here the extra term couples to four-velocity unit vectors $U_{\mu}U_{\nu}$ but not the metric, which, by using the definition of $T_{\mu\nu}$, suggests an effective energy density rather than pressure, given by
@@
Not surprisingly, we get analogous results to that found in the Brown picture. Here, we see the exchange of the derived quantity between $p$ in Brown’s picture and $\rho$ in Schutz’s picture led to the exchange of their effective role in the field equations. The expression for $\rho_{\rm eff}$ agrees with the on-shell expression in the argument presented in Sec. 2.3. Of course, one has to be careful when changing the variables $n$ and $\upmu$; however, it should be clear at this point that the trace coupling can only yield a contribution that affects the energy density or the pressure.
@@ 3.2 Equations of motion
$n\equiv\frac{\partial p}{\partial\upmu}+\frac{f_{T}}{2\kappa}\frac{\partial T}{\partial\upmu}\,;$ (3.19)
then we find from Eq. (3.17) that $\partial_{\nu}\big(nU^{\nu}\big)=0$. Using this relation and the product rule in Eqs. (3.15) and (3.18) allows us to write the equations of motion as
@@
The equations of motion (3.20)-(3.23) now match the expected result found by Schutz [12] with the new definition of $n$ in Eq. (3.19). We see, however, that the entropy variation $\delta s$ leads to an additional term $(2\kappa)^{-1}f_{T}\partial T/\partial s$ which is a result of the nonminimal coupling between curvature and the matter trace. In fact, even the factor $n$ in Eq. (3.23) which ensures the conservation of the particle number now depends on the nonminimal coupling between $R$ and $T$. Those effects will be discussed in more detail in Secs. 3.4 and 4.
@@ 3.3 On-shell fluid Lagrangian
$S_{\textrm{fluid}}=\int p(\mu,s)\sqrt{-g}\,d^{4}x\qquad\textrm{off shell}\,$, (3.25)
@@
$S_{\textrm{fluid}}$ $=-\int\rho_{\textrm{eff}}\sqrt{-g}\,d^{4}x\qquad\textrm{on shell}\,.$ (3.31)
Therefore, on shell, we do get back the expected expression after utilizing the equations of motion. This not only ensures the consistency of our results so far, but also strengthens the argument that one indeed has to be careful when working with the on-shell expression of the Lagrangian as it only holds up when the equations of motion are considered. However, when applying a variation principle, one must resort to the off-shell Lagrangian to get the full expression Eq. (3.10).
@@ 3.5 Conservation equations via field equations
We again check the consistency of our approach by first deriving the conservation equations from the field equations in this section and then deriving them once more from the fluid equation in the following section.
@@
where the contracted derivative of the entropy term is eliminated by the equation of motion (3.22). By substituting the expression for $T_{\mu\nu}^{\rm eff}$ from Eq. (3.12) we get
@@
Using the identities $h^{\nu}_{\sigma}h_{\mu\nu}=h^{\nu}_{\sigma}g_{\mu\nu}$ and $h^{\nu}_{\sigma}\nabla^{\mu}h_{\mu\nu}=h^{\nu}_{\sigma}U_{\mu}\nabla^{\mu}U_{\nu}$ allows us to combine the first two terms. After multiplying by $2\kappa$, the expression can be written in a form similar to that in Eq. (2.51):
@@
We have obtained the two expressions for the conservation equation parallel to the flow [Eq. (3.39)] and perpendicular to the flow [Eq. (3.42)]. Now, we attempt to rederive the same expressions using the fluid equation without resorting to the field equations.
@@ 3.6 Conservation equations via fluid equations
$\nabla^{\mu}T_{\mu\nu}$ $=\nabla^{\mu}\Big(\upmu\frac{\partial p}{\partial\upmu}U_{\mu}U_{\nu}\Big)+g_{\mu\nu}\nabla^{\mu}p\,$, (3.43)
@@
By expanding the derivative of the pressure and using Eq. (3.22),
@@
which matches the result from the field equations in Eq. (3.39), verifying our previous result that in this picture, the parallel contraction of the energy-momentum divergence does not vanish.
Now, we contract with $h^{\nu}_{\sigma}$ to check the conservation equations in the perpendicular direction:
@@
The term in round brackets vanishes [check Eq. (A.44) in Appendix A.5], leaving us with
@@
With that we have strong confidence that the correct approach to vary the fluid action is to use the full off-shell fluid Lagrangian. We showed that this can be done in both Brown’s approach with Lagrange multipliers or Schutz’s approach based on velocity potentials. The results in both cases are consistent and lead to the same conclusions.
3.7 Effective equations and dark sector interactions
We reexamine the mapping of our results to dark sector interactions. As before, we begin with $f(R,T)=R+\bar{f}(R,T)$ and find the field equation
$G_{\mu\nu}=\kappa T_{\mu\nu}+\frac{1}{2}\bar{f}_{T}\upmu\frac{\partial T}{\partial\upmu}U_{\mu}U_{\nu}+\frac{1}{2}\bar{f}g_{\mu\nu}-\big(R_{\mu\nu}+g_{\mu\nu}\Box-\nabla_{\mu}\nabla_{\nu}\big)\bar{f}_{R}$ (3.49)
and now follow a slightly different approach to before. The right-hand side contains an effective energy density term, unlike the Brown approach where the extra term contributed to the pressure. This means that the Schutz approach yields an extra contribution which is pressureless. Let us therefore write
$G_{\mu\nu}$ $=\kappa(T^{\rm eff}_{\mu\nu}+\hat{T}_{\mu\nu}^{\bar{f}})\,$, (3.50)
where $\hat{T}_{\mu\nu}^{\bar{f}}$ contains all other terms on the right-hand side of Eq. (3.49). Here, the effective energy-momentum tensor is as in Eq. (3.12), which means it contains the standard fluid plus a dark matter contribution. This formulation allows us to conclude that
$\nabla^{\mu}T^{\rm eff}_{\mu\nu}=-\nabla^{\mu}T_{\mu\nu}^{\bar{f}}=-\frac{1}{2\kappa}\bar{f}_{T}\nabla_{\nu}T=:\mathcal{Q}_{\nu}\,.$ (3.51)
It would now be natural to interpret this as follows: The effective energy-momentum tensor, which contains the fluid and a dark matter contribution, is not conserved. It leads to an energy transfer to $T_{\mu\nu}^{\bar{f}}$ which we would then interpret as the dark energy component.
The identification of the various terms with either dark matter or dark energy is not unique as we have mentioned before. However, using Schutz’s approach, one arrives at an effective energy density which lends itself to be interpreted as dark matter. In the other approach, see Sec. 2.7, this identification is not as direct. We would therefore conclude that the fluid Lagrangian based on the pressure; this means the approach by Schutz is preferable for a nonminimally coupled dark matter model.